OpenStudy (anonymous):

Use addition or substitution to find the value of x for this set of equations. y = x - 2 7y - 4x = 10

OpenStudy (anonymous):

keep the top as is, however the bottom you want to move the 7y over to the right behind the equal sign and -4x becomes positive 4x and move it over to the left. put 10 in the spot -4x was. This would cause the negative to be a positive sign making 10 a positive. your problem should now look like this: y=x-2 4x+10=7y Now you will want to substitute your y varable in the second equation with the y variable in the first equation. Now your answer should look like this: 4x+10=7(x-2) 4x+10=7x-14 now subtract both sides by -7 -3x+10=-14 is what youy should get now subtract -10 on both sides now you get -3x=-24 divde by both sides x=8 is your answer for x now plug in 8 into the first equation where x is the variable and you get: y=8-2 y=6 so now you have y =6 as your answer for y so your solutions are: {(8,6)}

OpenStudy (anonymous):

Substitution would be the best. You would want to "substitute the 1st equation into the second, and then rewrite it. So, it would look like this: 7(x-2)-4x=10 Next, use the distributive property: 7x-14-4x=10 Simply Simplify: 3x-14=10 Solve it (almost): 3x-14=10 +14 +14 Ok... 3x=24 Fianlly: X=8 So then after we find the X value, we plug this in to the Y= equation to get our Y value: y=x-2 y=(8)-2 Y=6 X=8 coordinates: [x,y] [8,6] Thank You. ~CHR